OpenStudy (anonymous):

Simplify. cuberoot(125n^12)

OpenStudy (anonymous):

\[\sqrt[3]{125n^{12}}\]

OpenStudy (anonymous):

the a should be a 3

OpenStudy (anonymous):

What's another way you could write this using a fractional exponent?

OpenStudy (anonymous):

i have no clue.

OpenStudy (anonymous):

How about this:\[(125n^{12})^\frac{1}{3} \]

OpenStudy (anonymous):

Now using rules of exponents, what should you get for n^12 raised to the 1/3?

OpenStudy (anonymous):

And for 125, actually.

OpenStudy (anonymous):

The rule for exponents is that you multiply, so now you have \[125^\frac{1}{3}*(n^{12})^\frac{1}{3}\]

OpenStudy (anonymous):

Oh ok, so then i would have 5n^4?

OpenStudy (anonymous):

Bingo! :)

OpenStudy (anonymous):

Awesome, thanks!

OpenStudy (anonymous):

Happy to help.